# Jarník-Besicovitch type theorems for semisimple algebraic groups

### v1

@@ Abstract.
In this note, we initiate a study on Jarník-Besicovitch type theorems for semisimple algebraic groups from the representation-theoretic point of view. Let $\rho:\mathbf{G}\to\operatorname{\mathbf{GL}}(V)$ be an irreducible $\mathbb{Q}$-rational representation of a connected semisimple $\mathbb{Q}$-algebraic group $\mathbf{G}$ on a complex vector space $V$ and $\{a_{t}\}_{t\in\mathbb{R}}$ a one-parameter subgroup in a $\mathbb{Q}$-split torus in $\mathbf{G}$. We define a subset $S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})$ of Diophantine elements of type $\tau$ in $\mathbf{G}(\mathbb{R})$ in terms of the representation $\rho$ and the subgroup $\{a_{t}\}_{t\in\mathbb{R}}$, and prove formulas for the Hausdorff dimension of the complement of $S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})$. As corollaries, we deduce several Jarník-Besicovitch type theorems.
@@ 1. Introduction and main results
1.1. Introduction
In this note, we initiate a study on the Jarník-Besicovitch theorem in the metric theory of Diophantine approximation. Recall that a real number $x\in\mathbb{R}$ is Diophantine of type $\kappa>0$ if there exists a constant $C>0$ such that
$\left|nx-m\right|\geq\frac{C}{n^{\kappa}}\;\left(\forall\frac{m}{n}\in\mathbb{Q}\right).$
Then necessarily $x\in\mathbb{R}\setminus\mathbb{Q}$ and the Dirichlet theorem implies that $\kappa\geq 1$. Denote by $S_{\kappa}$ the set of all Diophantine numbers of type $\kappa$ and $S_{\kappa}^{c}$ its complement. Then one can deduce from the Khintchine’s theorem on metric Diophantine approximation that $S_{\kappa}^{c}$ is a set of full measure if $\kappa=1$, and is null if $\kappa>1$. A refined result due to Jarník and Besicovitch [4, 22] states further that the Hausdorff dimension of $S_{\kappa}^{c}$ is equal to $2/(\kappa+1)$ ($\kappa>1$).
From the viewpoint of dynamical systems, the Jarník-Besicovitch theorem can be reformulated as a result of shrinking target problem, which was first observed by Hill and Velani [20]. Let $f:X\to X$ be a map on a metric space $X$ with a measure $\mu$. Generally speaking, in the shrinking target problem, one studies the set $S$ of points in $X$ whose trajectories under $f$ approach to a shrinking target infinitely often with certain rate, and seeks to establish results about the size (the measure or the Hausdorff dimension) of the set $S$. The measure theoretic version of the shrinking target problem is usually related to the Khintchine’s theorem [26, 41] while the Hausdorff dimension version is linked with the Jarník-Besicovitch theorem [20].
In homogeneous dynamics, various Diophantine sets can also be studied by the dynamical and ergodic properties of group actions on homogeneous spaces via Dani’s correspondence [14, 26]. Specifically, let $X_{m+n}=\operatorname{\mathbf{SL}}_{m+n}(\mathbb{R})/\operatorname{\mathbf{SL}}_{m+n}(\mathbb{Z})$,
@@ 1.1. Introduction
$a_{t}=\operatorname{diag}(e^{t/m},\dots,e^{t/m},e^{-t/n},\dots,e^{-t/n})\quad(t\in\mathbb{R})$
[…]
$g\operatorname{\mathbf{SL}}_{m+n}(\mathbb{Z})\to g\cdot\mathbb{Z}^{m+n}\quad(g\in\operatorname{\mathbf{SL}}_{m+n}(\mathbb{R})).$
Let $\delta:X_{m+n}\to\mathbb{R}_{+}$ be the systole function on $X_{m+n}$ defined by
$\delta(\Lambda)=\inf_{v\in\Lambda\setminus\{0\}}\|v\|$
which is also called the first minimum of the lattice $\Lambda\in X_{m+n}$. For any $m\times n$ matrix $A$, denote by
$u_{A}:=\left(\begin{array}[]{ccc}I_{m}&A\\ O&I_{n}\end{array}\right)\in N_{+}.$
Then it is well-known via Mahler’s compactness criterion and Dani’s correspondence [14, 26] that $A$ is badly approximable if and only if the orbit $\{a_{t}u_{A}\cdot\mathbb{Z}^{m+n}:t\geq 0\}$ is bounded in $X_{m+n}$, i.e.
$\delta(a_{t}u_{A}\cdot\mathbb{Z}^{m+n})\geq c\quad(t\geq 0)$
for some constant $c>0$; singular if and only if $\{a_{t}u_{A}\cdot\mathbb{Z}^{m+n}:t\geq 0\}$ diverges i.e.
$\delta(a_{t}u_{A}\cdot\mathbb{Z}^{m+n})\to 0\quad(\textup{as }t\to\infty);$
and very well approximable if and only if
$\limsup_{t\to\infty}\frac{-\log(\delta(a_{t}u_{A}\cdot\mathbb{Z}^{m+n}))}{t}>0.$
We say that an $m\times n$ matrix $A$ is Diophantine of type $\kappa$ if there exists a constant $C>0$ such that
$\|Aq-p\|\geq C\|q\|^{-\kappa}$
for any $q\in\mathbb{Z}^{n}\setminus\{0\}$ and $p\in\mathbb{Z}^{m}$. Then necessarily $\kappa\geq n/m$, and again using Dani’s correspondence, one can deduce that $A$ is Diophantine of type $\kappa$ if and only if there exists a constant $C>0$ such that
$\delta(a_{t}\cdot(u_{A}\cdot\mathbb{Z}^{m+n}))\geq Ce^{-\tau t}\;(\forall t>0)$
where
$\tau=\frac{m\kappa-n}{mn(\kappa+1)}.$
The set $S_{\kappa}$ of all Diophantine numbers of type $\kappa$ defined at the beginning of the introduction then corresponds to the case $m=n=1$ here. In particular, one can rephrase the Jarník-Besicovitch theorem for the homogeneous space $X_{m+n}=X_{2}$ $(m=n=1)$, and get that for any parameter $10\leq\tau<$ 1, the Hausdorff dimension of the complement of the set
$\{u_{A}\in N_{+}:\textup{there exists $C>0$ such that }\delta(a_{t}\cdot(u_{A}\cdot\mathbb{Z}^{2}))\geq Ce^{-\tau t}\;(\forall t>0)\}$
is equal to $1-\tau$. Using the decomposition of $X_{2}$ into the stable, central and unstable submanifolds of the flow $\{a_{t}\}_{t\in\mathbb{R}}$ on $X_{2}$, one can then obtain the following equivalent statement: the Hausdorff dimension of the complement of the set
$\{p\in X_{2}:\textup{there exists $C>0$ such that }\delta(a_{t}\cdot p)\geq Ce^{-\tau t}\;(\forall t>0)\}$
is equal to $3-\tau$. In general, the matrix version of the Jarnik-Besicovitch theorem has also been established [9, 15], and it is equivalent to the statement that for any $0\leq\tau<1/n$, the Hausdorff dimension of the complement of the set
$\{p\in X_{m+n}:\textup{there exists $C>0$ such that }\delta(a_{t}\cdot p)\geq Ce^{-\tau t}\;(\forall t>0)\}$
is equal to $mn(1-\tau)+m^{2}+n^{2}+mn-1$. We remark that one can further consider the case where $\{a_{t}\}_{t\in\mathbb{R}}$ is a generic one-parameter diagonal subgroup in $\operatorname{\mathbf{SL}}_{m+n}(\mathbb{R})$ and study the Hausdorff dimension of the complement of the set
$\{p\in X_{m+n}:\textup{there exists $C>0$ such that }\delta(a_{t}\cdot p)\geq Ce^{-\tau t}\;(\forall t>0)\}.$
This type of question leads to a general weighted multidimensional Jarník-Besicovitch theorem on metric Diophantine approximation, which is one of the main topics we pursue in this note (See Theorem 1.6).
[…]
@@
Let $\{a_{t}\}_{t\in\mathbb{R}}$ be a one-parameter Ad-diagonalizable subgroup in $\mathbf{G}(\mathbb{R})$. Then the main object we would like to study is the Hausdorff dimension of the complement of the set
$\{g\in\mathbf{G}(\mathbb{R}):\delta(\rho(a_{t}\cdot g)\cdot\mathbb{Z}^{d})\geq Ce^{-\tau t}\textup{ for some }C>0\}.$
Clearly the discussions above about the Jarnik-Besicovitch theorem with the homogeneous flow $a_{t}=\operatorname{diag}(e^{t/m},\dots,e^{t/m},e^{-t/n},\dots,e^{-t/n})$ on $X_{m+n}$ can be reduced to this setting if we let $\mathbf{G}=\operatorname{\mathbf{SL}}_{m+n}$ and $\rho$ the standard representation of $\operatorname{\mathbf{SL}}_{m+n}$ on $\mathbb{C}^{m+n}$. Moreover, we will see later (Theorem 1.8) that by considering representations of algebraic groups, one will be able to establish some Jarnik-Besicovitch type theorems for certain algebraic subvarieties in an affine space. We remark that the question we study in this note starts from the sparse equidistribution problem, and a related shrinking target problem plays an important role. For more details, one can read [17, 46, 47, 48].
@@ 1.2. Main results
In the rest of this note, we address the question above under the following assumption. We assume that $\mathbf{G}$ is a connected semisimple $\mathbb{Q}$-algebraic group (in the Zariski topology) and $\mathbf{T}$ is a maximal $\mathbb{Q}$-split torus in $\mathbf{G}$. Let $\{a_{t}\}_{t\in\mathbb{R}}$ be a one-parameter subgroup in $\mathbf{T}(\mathbb{R})$, and $\rho$ a finite-dimensional irreducible representation of $\mathbf{G}$ defined over $\mathbb{Q}$ on a complex vector space $V$ with $\dim\ker\rho=0$. Let $d=\dim V$ and we may identify $V\cong\mathbb{C}^{d}$ so that $\mathbb{Z}^{d}\subset\mathbb{C}^{d}$ is compatible with the $\mathbb{Q}$-structure in $V$.
@@
An element $g\in\mathbf{G}(\mathbb{R})$ is Diophantine of type $\tau\geq 0$ if there exists a constant $C>0$ such that
$\delta(\rho(a_{t}\cdot g)\mathbb{Z}^{d})\geq Ce^{-\tau t}\textup{ for any }t>0$
where $\delta$ is the first minimum function. We denote by $S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})$ the set of Diophantine elements of type $\tau$ in $\mathbf{G}(\mathbb{R})$ and $S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})^{c}$ its complement in $\mathbf{G}(\mathbb{R})$. If the representation $\rho$ and the one-parameter subgroup $\{a_{t}\}_{t\in\mathbb{R}}$ are clearly stated in the contexts, we will simply write $S_{\tau}$ instead of $S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})$.
@@
If $\mathbf{G}$ is $\mathbb{Q}$-anisotropic, then $\mathbf{T}=\{e\}$. So in the following, we assume that $\mathbf{G}$ is $\mathbb{Q}$-isotropic.
To state the first main theorem about $\dim_{H}S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})^{c}$, we need to introduce some notation. We choose a minimal parabolic $\mathbb{Q}$-subgroup $\mathbf{P}_{0}$ in $\mathbf{G}$ containing $\mathbf{T}$. Then $\mathbf{P}_{0}$ and $\mathbf{T}$ defines a root system ($\Phi,\;\Phi^{+},\;\Delta$) where $\Phi$ is the set of $\mathbb{Q}$-roots relative to $\mathbf{T}$, $\Phi^{+}$ is the set of positive $\mathbb{Q}$-roots determined by $\mathbf{P}_{0}$ and $\Delta$ is the set of simple $\mathbb{Q}$-roots in $\Phi^{+}$. Let $\overline{\mathbf{P}}_{0}$ be the opposite minimal parabolic $\mathbb{Q}$-subgroup of $\mathbf{P}_{0}$ defined by $\Phi\setminus\Phi^{+}$. Without loss of generality, we may assume that the stable horospherical subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$ is contained in the unipotent radical $R_{u}(\mathbf{P}_{0})$ of $\mathbf{P}_{0}$ and its unstable horospherical subgroup is contained in the unipotent radical $R_{u}(\overline{\mathbf{P}}_{0})$ of $\overline{\mathbf{P}}_{0}$. One can write the space $V$ in the representation $\rho$ as a direct sum of weight spaces with respect to the action of $\mathbf{T}$
@@
By the structure theory of irreducible representations of complex semisimple groups and semisimple Lie algebras, there is a highest weight $\beta_{0}$ among the weights $\beta$’s (where the order is determined by the minimal parabolic $\mathbb{Q}$-subgroup $\mathbf{P}_{0}$) and we denote by $V_{\beta_{0}}$ its corresponding weight space. (See §2 for more details). The stabilizer of the weight space $V_{\beta_{0}}$ in $\mathbf{G}$ is a parabolic $\mathbb{Q}$-subgroup $\mathbf{P}_{\beta_{0}}$ containing $\mathbf{P}_{0}$, and its unipotent radical is denoted by $R_{u}(\mathbf{P}_{\beta_{0}})$. The opposite parabolic $\mathbb{Q}$-subgroup of $\mathbf{P}_{\beta_{0}}$ containing $\overline{\mathbf{P}}_{0}$ and its unipotent radical are denoted by $\overline{\mathbf{P}}_{\beta_{0}}$ and $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$ respectively.
In the following, if an algebraic $\mathbb{Q}$-subgroup $\mathbf{F}\subset\mathbf{G}$ is normalized by $\mathbf{T}$, then we write $\Phi(\mathbf{F})$ for the set of $\mathbb{Q}$-roots in $\mathbf{F}$ relative to $\mathbf{T}$ while the summation $\sum_{\alpha\in\Phi(\mathbf{F})}$ (or $\prod_{\alpha\in\Phi(\mathbf{F})}$) means that we take the sum (or product) all over $\alpha\in\Phi(\mathbf{F})$ counted with multiplicities (i.e., the dimensions of the corresponding $\mathbb{Q}$-root spaces in the Lie algebra of $\mathbf{F}$). For a $\mathbb{Q}$-root or a $\mathbb{Q}$-weight $\lambda$ in $\rho$ relative to $\mathbf{T}$ (which is a $\mathbb{Q}$-character of $\mathbf{T}$) , we will often consider it as a linear functional on the Lie algebra $\mathfrak{a}$ of $\mathbf{T}(\mathbb{R})$ and use the same symbol. In particular, we will write $\lambda(a_{t})$ ($t\in\mathbb{R}$) for the values of $\lambda$ (as a linear functional) on the Lie algebra of $\{a_{t}\}_{t\in\mathbb{R}}$ (so that $\lambda(a_{t})$ is linear on $t\in\mathbb{R}$). We denote by $\nu_{0}$ the $\mathbb{Q}$-root in $R_{u}(\overline{\mathbf{P}}_{0})$ relative to $\mathbf{T}$ such that
@@
Now we can state the first main theorem in this note.
@@
Let $\mathbf{G}$ be a connected semisimple algebraic group defined over $\mathbb{Q}$, $\mathbf{T}$ a maximal $\mathbb{Q}$-split torus in $\mathbf{G}$ and $\{a_{t}\}_{t\in\mathbb{R}}$ a one-parameter subgroup in $\mathbf{T}(\mathbb{R})$. Let $\rho$ be a finite-dimensional irreducible representation of $\mathbf{G}$ defined over $\mathbb{Q}$ on a complex vector space $V$ with $\dim\ker\rho=0$. Then
$\dim_{H}S_{\tau}(\rho,\{a_{t}\}_{t\in\mathbb{R}})^{c}\geq$ $\dim\mathbf{G}-\frac{\tau}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}\cdot\sum_{\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{\beta_{0}}))}\alpha(a_{1})$
for any $0\leq\tau<\beta_{0}(a_{-1})$.
To state the next theorem, we need to introduce another notation. Let $N(\mathbf{T})$ and $Z(\mathbf{T})$ be the normalizer and centralizer of $\mathbf{T}$ respectively. Then the Weyl group relative to $\mathbb{Q}$ is defined by

### v2

# A shrinking target problem in homogeneous spaces of semisimple algebraic groups

@@ Abstract.
In this paper, we study a shrinking target problem with target at infinity in a homogeneous space of a semisimple algebraic group from the representation-theoretic point of view. Let $\rho:\mathbf{G}\to\operatorname{\mathbf{GL}}(V)$ be an irreducible $\mathbb{Q}$-rational representation of a connected semisimple $\mathbb{Q}$-algebraic group $\mathbf{G}$ on a complex vector space $V$, $\{a_{t}\}_{t\in\mathbb{R}}$ a one-parameter subgroup in a $\mathbb{Q}$-split torus in $\mathbf{G}$ and $\psi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ a positive function on $\mathbb{R}_{+}$. We define a subset $S_{\rho}(\psi)$ of $\psi$-Diophantine elements in $\mathbf{G}(\mathbb{R})$ in terms of the representation $\rho$ and $\{a_{t}\}_{t\in\mathbb{R}}$, and prove formulas for the Hausdorff dimension of the complement of $S_{\rho}(\psi)$. We also discuss the connections of our results to Diophantine approximation on flag varieties and rational approximation to linear subspaces in Grassmann varieties.
Key words and phrases:
shrinking target problem, Dani correspondence, structures of irreducible representations of semisimple algebraic groups, reduction theory of arithmetic subgroups, mixing of a semisimple flow, asymptotic estimates of the number of rational points, Diophantine approximation on flag varieties
@@ 1. Introduction and main results
1.1. Introduction and main problem
In this paper, we study a shrinking target problem, the prototype of which was first proposed by Hill and Velani [39]. Let $f:X\to X$ (or $f_{t}:X\to X$) be a map (or a flow) on a metric space $X$ with a measure $\mu$. Generally speaking, in the shrinking target problem, one studies the set $S$ of points in $X$ whose orbits under $f$ (or $f_{t}$) hit a shrinking target infinitely often, and seeks to establish results about the size (the $\mu$-measure or the Hausdorff dimension) of $S$. The question about the measure of $S$, especially when $f_{t}$ is a flow on a homogeneous space, is closely related to Khintchine-type theorems in the metric theory of Diophantine approximation [47, 48, 35], and has been studied extensively in the past few decades (e.g. [2, 32, 45, 46, 50, 54, 77, 79, 53, 5]). In particular, the shrinking target problem of this type for a diagonalizable homogeneous flow $f_{t}$ has been addressed in full generality by Kleinbock and Margulis [50] where a null-conull law on $\mu(S)$ and a logarithm law are established. On the other hand, the study of the Hausdorff dimension of $S$ for a homogeneous flow $f_{t}$ (when $\mu(S)=0$) is a finer problem, and it is usually related to the Jarník-Besicovitch theorem in Diophantine approximation [6, 42]. As the analysis of the Hausdorff dimension of $S$ is more delicate and requires more information about the geometric structure of the space $X$ and the dynamics of the homogeneous flow $f_{t}$, only a few cases were known (which we will discuss below in this subsection). In this paper, we focus on this Hausdorff dimension version of shrinking target problem for homogeneous flows. (For discussions about other maps or flows, one may refer to e.g. [23, 40, 58, 70, 74, 81, 57].)
In [82, 25], we consider a shrinking target problem with target at infinity for the homogeneous flow $a_{t}:G/\Gamma\to G/\Gamma$ where $G/\Gamma$ is a finite-volume quotient of a rank-one simple Lie group $G$ [82] or $G/\Gamma=\operatorname{\mathbf{SL}}_{3}(\mathbb{R})/\operatorname{\mathbf{SL}}_{3}(\mathbb{Z})$ [25], and $\{a_{t}\}_{t\in\mathbb{R}}$ is a one-parameter diagonalizable subgroup in $G$. In particular, we obtain a formula for the Hausdorff dimension of the set $S$ as defined above, and establish a Jarník-Besicovitch theorem on Diophantine approximation in Heisenberg groups [82]. One may also refer to [38, 36, 37, 62] for related discussions.
This paper is a continuation of the works [82, 25], and we aim to generalize the main results in [82, 25] for diagonalizable homogeneous flows on finite-volume quotients of semisimple algebraic groups. We will see that our generalization has close connections to Diophantine approximation on flag varieties [56, 17] and rational approximation to linear subspaces in Grassmann varieties [71, 16], and it can imply certain Jarník-Besicovitch type theorems about well approximable subsets. Currently there are references [11, 20, 15, 75] which may be related to the topic of this paper with different emphases. In [11, 20], the main results are Jarník-Besicovitch theorems about Diophantine matrices, but with the help of a generalized Dani correspondence developed in [14, 50], these results can be reformulated as a shrinking target problem for the homogeneous flow $h_{t}:X_{m+n}\to X_{m+n}$ where $X_{m+n}=\operatorname{\mathbf{SL}}_{m+n}(\mathbb{R})/\operatorname{\mathbf{SL}}_{m+n}(\mathbb{Z})$ and
@@ 1.1. Introduction and main problem
$h_{t}=\operatorname{diag}(\underbrace{e^{t/m},\dots,e^{t/m}}_{m\textup{ times}},\underbrace{e^{-t/n},\dots,e^{-t/n}}_{n\textup{ times}})\quad(t\in\mathbb{R}).$
In [15], a variational principle is established to analyze $h_{t}$-orbits with various behaviors in $X_{m+n}$, which makes it possible to study the Hausdorff dimensions of a variety of Diophantine subsets in the space of $m\times n$ matrices. This variational principle is generalized in [75] for any diagonalizable flow on $X_{m+n}$ with respect to a nonstandard Hausdorff dimension. We remark that the results in [11, 20, 15, 75] mainly deal with diagonalizable flows on $X_{m+n}$.
Now we propose the main problem of this paper. Let $\mathbf{G}$ be a semisimple algebraic group defined over $\mathbb{Q}$ and $\rho:\mathbf{G}\to\operatorname{\mathbf{GL}}(V)$ a finite-dimensional irreducible representation of $\mathbf{G}$ over $\mathbb{Q}$ on a complex vector space $V$ with a $\mathbb{Q}$-structure. The vector space $V$ may be identified with $\mathbb{C}^{d}$ ($d=\dim_{\mathbb{C}}V$) equipped with a norm $\|\cdot\|$ so that $\mathbb{Z}^{d}\subset\mathbb{C}^{d}$ is compatible with the $\mathbb{Q}$-structure in $V$. For any discrete subgroup $\Lambda$ in $V$, define the first minimum of $\Lambda$ by
@@
Let $\{a_{t}\}_{t\in\mathbb{R}}$ be a one-parameter Ad-diagonalizable subgroup in $\mathbf{G}(\mathbb{R})$, $\mathbb{R}_{+}$ the set of positive real numbers and $\psi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ a positive function on $\mathbb{R}_{+}$. Then we aim to estimate the Hausdorff dimension of the complement of the set
$\{g\in\mathbf{G}(\mathbb{R}):\delta(\rho(a_{t}\cdot g)\cdot\mathbb{Z}^{d})\geq C\cdot\psi(t)\;(\forall t>0)\textup{ for some }C>0\}\subset\mathbf{G}(\mathbb{R})$
with respect to a standard Riemannian metric on $\mathbf{G}(\mathbb{R})$.
By Mahler’s compactness criterion, the main problem above is clearly a shrinking target problem with target at infinity for the homogeneous flow $a_{t}:\mathbf{G}(\mathbb{R})/\Gamma\to\mathbf{G}(\mathbb{R})/\Gamma$, where $\Gamma$ is an arithmetic lattice in $\mathbf{G}(\mathbb{R})$ preserving $\mathbb{Z}^{d}$. Note that the results in [11, 20] can be translated in the setting of this problem if we let $\mathbf{G}=\operatorname{\mathbf{SL}}_{m+n}$, $\rho$ the standard representation of $\operatorname{\mathbf{SL}}_{m+n}$ on $\mathbb{C}^{m+n}$ and $\{a_{t}\}_{t\in\mathbb{R}}=\{h_{t}\}_{t\in\mathbb{R}}$. We will see later in our theorems that the main problem also includes the cases in [82, 25].
@@ 1.2. Main results
In this paper, we address the main problem above under the following assumption. We assume that $\mathbf{G}$ is a connected semisimple algebraic group over $\mathbb{Q}$ (with respect to the Zariski topology) and $\mathbf{T}$ is a maximal $\mathbb{Q}$-split torus in $\mathbf{G}$. Let $\{a_{t}\}_{t\in\mathbb{R}}$ be a one-parameter subgroup in $\mathbf{T}(\mathbb{R})$, $\psi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ a positive function on $\mathbb{R}_{+}$ and $\rho$ a finite-dimensional irreducible representation of $\mathbf{G}$ defined over $\mathbb{Q}$ with $\dim\ker\rho=0$.
@@
An element $g\in\mathbf{G}(\mathbb{R})$ is called $\psi$-Diophantine if there exists a constant $C>0$ such that
$\delta(\rho(a_{t}\cdot g)\mathbb{Z}^{d})\geq C\cdot\psi(t)\textup{ for any }t>0$
where $\delta$ is the first minimum function. We denote by $S_{\rho}(\psi)$ the set of all $\psi$-Diophantine elements, and $S_{\rho}(\psi)^{c}$ its complement in $\mathbf{G}(\mathbb{R})$.
@@
In the rest of the paper, we assume that $\mathbf{G}$ is $\mathbb{Q}$-isotropic since otherwise $\mathbf{T}=\{e\}$. Note that for any $g\in\mathbf{G}(\mathbb{R})$, $\rho(a_{t}\cdot g)\cdot\mathbb{Z}^{d}$ is a unimodular lattice in the vector space of real points in $V$, and $\delta(\rho(a_{t}\cdot g)\cdot\mathbb{Z}^{d})$ is bounded above by a constant depending only on $\rho$. Due to this fact, we may also assume that the function $\psi$ is bounded since otherwise $S_{\rho}(\psi)=\emptyset$ by Definition 1.1. We denote by $\dim_{H}(S)$ the Hausdorff dimension of a subset $S$ in a smooth manifold $\mathcal{M}$ with respect to a Riemannian metric on $\mathcal{M}$.
To state the first two theorems about $\dim_{H}S_{\rho}(\psi)^{c}$, we need to introduce some notation. We choose a minimal parabolic $\mathbb{Q}$-subgroup $\mathbf{P}_{0}$ in $\mathbf{G}$ containing $\mathbf{T}$ with the Levi subgroup $Z(\mathbf{T})$ (the centralizer of $\mathbf{T}$ in $\mathbf{G}$). Then $\mathbf{P}_{0}$ and $\mathbf{T}$ define a root system ($\Phi,\;\Phi^{+},\;\Delta$) where $\Phi$ is the set of $\mathbb{Q}$-roots relative to $\mathbf{T}$, $\Phi^{+}$ is the set of positive $\mathbb{Q}$-roots determined by $\mathbf{P}_{0}$ and $\Delta$ is the set of simple $\mathbb{Q}$-roots in $\Phi^{+}$. Let $\overline{\mathbf{P}}_{0}$ be the opposite minimal parabolic $\mathbb{Q}$-subgroup of $\mathbf{P}_{0}$ defined by $\Phi\setminus\Phi^{+}$ with the same Levi subgroup $Z(\mathbf{T})$. Without loss of generality, we may assume that the stable horospherical subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$ is contained in the unipotent radical $R_{u}(\mathbf{P}_{0})$ of $\mathbf{P}_{0}$ and the unstable horospherical subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$ is contained in the unipotent radical $R_{u}(\overline{\mathbf{P}}_{0})$ of $\overline{\mathbf{P}}_{0}$. One can write the space $V$ in the representation $\rho$ as a direct sum of weight spaces with respect to the action of $\mathbf{T}$
@@
By the structure theory of irreducible representations of complex semisimple groups and semisimple Lie algebras, there is a highest weight $\beta_{0}$ among the weights $\beta$’s (where the order is determined by $\Phi^{+}$), and we denote by $V_{\beta_{0}}$ its corresponding weight space. (See §2 for more details). The stabilizer of the weight space $V_{\beta_{0}}$ in $\mathbf{G}$ is a parabolic $\mathbb{Q}$-subgroup $\mathbf{P}_{\beta_{0}}$ containing $\mathbf{P}_{0}$, and its unipotent radical is denoted by $R_{u}(\mathbf{P}_{\beta_{0}})$. It is known that there exists a $\mathbb{Q}$-torus $\mathbf{T}_{\beta_{0}}$ in $\mathbf{T}$ such that $Z(\mathbf{T}_{\beta_{0}})$ (the centralizer of $\mathbf{T}_{\beta_{0}}$ in $\mathbf{G}$) is a Levi subgroup of $\mathbf{P}_{\beta_{0}}$. We denote by $\overline{\mathbf{P}}_{\beta_{0}}$ the opposite parabolic $\mathbb{Q}$-subgroup of $\mathbf{P}_{\beta_{0}}$ containing $\overline{\mathbf{P}}_{0}$ with the same Levi subgroup $Z(\mathbf{T}_{\beta_{0}})$. The unipotent radical of $\overline{\mathbf{P}}_{\beta_{0}}$ is denoted by $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$.
In the following, if an algebraic $\mathbb{Q}$-subgroup $\mathbf{F}\subset\mathbf{G}$ is normalized by $\mathbf{T}$, then we write $\Phi(\mathbf{F})$ for the set of $\mathbb{Q}$-roots in $\mathbf{F}$ relative to $\mathbf{T}$, while the symbol $\sum_{\alpha\in\Phi(\mathbf{F})}$ (or $\prod_{\alpha\in\Phi(\mathbf{F})}$) stands for the sum (or product) over all the $\mathbb{Q}$-roots $\alpha\in\Phi(\mathbf{F})$ counted with multiplicities (i.e., the dimensions of the corresponding $\mathbb{Q}$-root spaces associated to $\alpha\in\Phi(\mathbf{F})$ in the Lie algebra of $\mathbf{F}$). For a $\mathbb{Q}$-root $\lambda$ in $\mathbf{G}$ or a $\mathbb{Q}$-weight $\lambda$ in $\rho$ relative to $\mathbf{T}$ (which is a $\mathbb{Q}$-character of $\mathbf{T}$), we will often consider it as a linear functional on the Lie algebra $\mathfrak{a}$ of $\mathbf{T}(\mathbb{R})$ and use the same symbol. In particular, we will write $\lambda(a_{t})$ ($t\in\mathbb{R}$) for the values of $\lambda$ (as a linear functional) on the Lie algebra of $\{a_{t}\}_{t\in\mathbb{R}}$ (so that $\lambda(a_{t})$ is linear on the parameter $t\in\mathbb{R}$). We denote by $\nu_{0}$ any $\mathbb{Q}$-root relative to $\mathbf{T}$ in $R_{u}(\overline{\mathbf{P}}_{0})$ satisfying the condition that
@@
Since the unstable horospherical subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$ is contained in $R_{u}(\overline{\mathbf{P}}_{0})$, the value $\nu_{0}(a_{1})>0$.
Now we can state our first two theorems. Here for any $\psi:\mathbb{R}_{+}\to\mathbb{R}_{+}$, we define
$\tau(\psi):=\liminf_{t\to\infty}\left(-\frac{\ln(\psi(t))}{t}\right)$
which will be an important quantity in the formula of the Hausdorff dimension of $S_{\rho}(\psi)^{c}$. We prove that the effective range of $\tau(\psi)$ is $0\leq\tau(\psi)\leq\beta_{0}(a_{-1})$.
@@
We have $\tau(\psi)\geq 0$ and $\beta_{0}(a_{-1})>0$. If $\tau(\psi)>\beta_{0}(a_{-1})$, then $S_{\rho}(\psi)^{c}=\emptyset$.
Theorem 1.4.
Let $\mathbf{G}$ be a connected semisimple algebraic group defined over $\mathbb{Q}$, $\mathbf{T}$ a maximal $\mathbb{Q}$-split torus in $\mathbf{G}$ and $\{a_{t}\}_{t\in\mathbb{R}}$ a one-parameter subgroup in $\mathbf{T}(\mathbb{R})$. Let $\rho$ be a finite-dimensional irreducible representation of $\mathbf{G}$ defined over $\mathbb{Q}$ on a complex vector space $V$ with $\dim\ker\rho=0$, and $\psi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ a function on $\mathbb{R}_{+}$. Then
$\dim_{H}S_{\rho}(\psi)^{c}\geq$ $\dim\mathbf{G}-\frac{\tau(\psi)}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}\cdot\sum_{\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{\beta_{0}}))}\alpha(a_{1})$
for any $0\leq\tau(\psi)<\beta_{0}(a_{-1})$. In the case $\tau(\psi)=\beta_{0}(a_{-1})$ we have
  1. (1)
    If $\psi(t)\cdot e^{\beta_{0}(a_{-1})t}$ is bounded, then $S_{\rho}(\psi)^{c}=\emptyset$.
  2. (2)
    If $\psi(t)\cdot e^{\beta_{0}(a_{-1})t}$ is unbounded, then $S_{\rho}(\psi)^{c}\neq\emptyset$ and
    $\dim_{H}S_{\rho}(\psi)^{c}\geq\dim\mathbf{G}-\frac{\tau(\psi)}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}\cdot\sum_{\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{\beta_{0}}))}\alpha(a_{1}).$
To state the next theorem, we need to introduce another notation. Let $N(\mathbf{T})$ and $Z(\mathbf{T})$ be the normalizer and centralizer of $\mathbf{T}$ in $\mathbf{G}$ respectively. Then the Weyl group relative to $\mathbb{Q}$ is defined by